Please help me with this problem.
Find the real numberx x and y
23+yi=17-x+3i

Answer by RAY100(1637)

23 + yi = 17-x +3i
.
separate real and imaginary terms
.
( 23) +{ yi} = (17-x) + { 3i}
.
set real = real,,,,,imaginary = imaginary
.
23 = 17 -x
x= (-6)
.
yi = 3 i
y=3
.
check by subst in origonal
23 +3i = 17 - (-6) +3i =23 +3i ,,,,,ok